J.Tamata shares with you the following information regarding his investment portfolio: Investment in four shares CFL, ATH, FTV & KFL. Total Portfolio is worth $100,000. Investment in ATH is half the size of CFL. Investment in FTV is half the size of ATH. Investment in KFL is half the size of FTV. Weighted beta of CFL and KFL are 0.35 and 0.55 respectively. Market is 0.15 more risky and 0.15 less risky for FTV and ATH shares respectively.

Calculate

a. The total dollar value and weight of each share in J. Tamata portfolio.

b. Beta of CFL.

c. Beta of KFL.

d. Weighted beta of ATH

e. Weighted beta of FTV

f. The portfolio weighted beta.

g.Tamata adds to his current portfolio shares of RBG worth of $10,000. Re calculate the portfolios new weighted beta and briefly explain the impact RBG share purchases has on the risk structure of J. Tamata portfolio? RBG shares has a beta of 2.3


1
Expert's answer
2021-05-30T14:14:50-0400

a) Investment in KFL =x=6.67%=6670

Investment in FTV=2x=13.34%=13340

Investment in ATH=4x=26.68%=26680

Investment in CFL=8x=53.36%=53360

"x+2x+4x+8x=100\%"

"x=6.67\%"

b)"b*53.36\%=0.35" "b=0.6559"

c)"b*6.67\%=0.55" "b=8.2459"

d)"1.15*26.68\%=0.3068"

e)"0.85*13.34\%=0.1134"

f)"0.35+0.55+0.31+0.11=1.32"

g) KFL "=6670\/110000=0.0606"

FTV"=13340\/110000=0.1213"

ATH"=26680\/110000=0.2425"

CFL"=53360\/110000=0.4851"

RBG"=10000\/110000=0.0909"

the portfolios new weighted beta:"8.2459*0.0606+0.85*0.1213+1.15*0.2425+0.6559*0.4851+2.3*0.0909=1.41"

Portfolio risk has increased as the new shareholding is a riskier than the market